### v1

# Editing the Mind of Giants: An In-Depth Exploration of Pitfalls of Knowledge Editing in Large Language Models

Abstract
Knowledge editing is a rising technique for efficiently updating factual knowledge in Large Language Models (LLMs) with minimal alteration of parameters. However, recent studies have identified concerning side effects, such as knowledge distortion and the deterioration of general abilities, that have emerged after editing. This survey presents a comprehensive study of these side effects, providing a unified view of the challenges associated with knowledge editing in LLMs. We discuss related works and summarize potential research directions to overcome these limitations. Our work highlights the limitations of current knowledge editing methods, emphasizing the need for deeper understanding of inner knowledge structures of LLMs and improved knowledge editing methods. To foster future research, we have released the complementary materials such as paper collection publicly [^1] https://github.com/MiuLab/EditLLM-Survey. **footnotetext: Equal contribution.
Editing the Mind of Giants: An In-Depth Exploration of Pitfalls of Knowledge Editing in Large Language Models
Cheng-Hsun Hsueh Paul Kuo-Ming Huang Tzu-Han Lin Che-Wei Liao Hung-Chieh Fang Chao-Wei Huang Yun-Nung Chen National Taiwan University, Taipei, Taiwan {r12922059, b08902072, b08902126, r09922a25}@csie.ntu.edu.tw {b09902106, f07922069}@csie.ntu.edu.tw y.v.chen@ieee.org
1 Introduction
Recent advancements in large language models (LLMs) have significantly improved Natural Language Processing (NLP; Brown et al., 2020). LLMs can now understand and generate language at a human-like level, establishing them as milestones in generative AI. This proficiency is due to their ability to extract knowledge from extensive text corpora. However, the mechanisms of knowledge storage in LLMs are not well understood, leading to potential issues in real-world applications. For instance, deploying LLMs as chatbots raises concerns about the reliability of their generated content due to the unclear mechanisms of knowledge storage and potential inaccuracies.
Figure 1: An overview of pitfalls in current knowledge editing methods. The subsequent sections dive into three key challenges: generalization issues (Section 3.1), localization issues (Section 3.2), and deterioration of general LLM abilities (Section 3.3).
To address these issues, researchers have explored various methods. Traditional approaches like fine-tuning, continual learning, and retraining are either computationally expensive or detrimental to the LLMs’ overall performance. Recently, knowledge editing has emerged as a promising alternative, offering minimal computational costs and fewer alterations (Cao et al., 2021; Dai et al., 2022; Meng et al., 2022, 2023; Dong et al., 2022; Mitchell et al., 2022a, b; Hartvigsen et al., 2023; Huang et al., 2023; Yu et al., 2024; Zheng et al., 2023; Li et al., 2024a; Tan et al., 2024; Gupta et al., 2024b; Wang et al., 2024). This approach allows for precise adjustments to LLMs, enhancing their practicality and reliability in real-world applications.
Knowledge editing can be divided into two main categories: parameter-modifying and parameter-preserving. Both aim to refine LLM knowledge efficiently while avoiding the drawbacks of previous tuning methods (Yao et al., 2023a). Parameter-modifying methods, including meta-learning (Cao et al., 2021; Mitchell et al., 2022a; Tan et al., 2024) and locate-then-edit techniques (Dai et al., 2022; Meng et al., 2022, 2023; Li et al., 2024a; Gupta et al., 2024b), strive to update a targeted set of model parameters while making effective edits. Parameter-preserving methods introduce external components, like knowledge bases (Mitchell et al., 2022b; Yu et al., 2024) or extra model parameters (Dong et al., 2022; Huang et al., 2023; Hartvigsen et al., 2023; Yu et al., 2024) to maintain the integrity of pre-trained LLMs while updating their knowledge.
Despite the success of knowledge editing, challenges remain. Knowledge editing can have unintended side effects, potentially damaging the general abilities and intrinsic structures of LLMs. Previous research has mainly focused on performance improvements and innovations within knowledge editing methods, with limited attention to potential drawbacks. Consequently, this survey aims to provide a holistic view of current issues in the knowledge editing paradigm and encourage further investigations into the pitfalls and intrinsic knowledge structures of LLMs. A brief overview of the discussed pitfalls is shown in Figure  1.
This survey is organized as follows: Section 2 introduces knowledge editing and its methods. Section 3 discusses current challenges, benchmarks, and methodologies. In Section 4, we present experimental results evaluating different editing methods. Finally, Section 5 explores related studies and future research directions. We summarize our contributions as follows:
  1. 1.
    We are the first to provide a comprehensive analysis of the side effects associated with existing knowledge editing techniques.
  2. 2.
    We systematically organized previous research and conducted experiments to benchmark the side effects of knowledge editing, providing a unified perspective on this issue.
  3. 3.
    We discussed related studies and potential research directions to address existing challenges, encouraging further exploration and understanding in this field.
Figure 2: Illustration of properties that knowledge editing methods should satisfy. An ideal knowledge editing method should be reliable, generalizable to relevant queries, and should not alter the outputs of irrelevant queries.
Figure 3: Illustration of the two categories of model editing methods in transformer-based large language models, which includes parameter-modifying (meta-learning and locate-and-edit) and parameter-preserving (additional parameters, external memory, in-context learning, and decoding) methods. MHSA and FFN stand for multi-head self-attention and feed-forward network, respectively.
Category Strategy Method
Parameter- modifying Meta- learning Knowledge Editor (Cao et al., 2021)
MEND (Mitchell et al., 2022a)
MALMEN (Tan et al., 2024)
Locating and editing Knowledge Neuron (Dai et al., 2022)
ROME (Meng et al., 2022)
MEMIT (Meng et al., 2023)
PMET (Li et al., 2024a)
EMMET (Gupta et al., 2024b)
Parameter- preserving Additional parameters CaliNET (Dong et al., 2022)
T-Patcher (Huang et al., 2023)
GRACE (Hartvigsen et al., 2023)
MELO (Yu et al., 2024)
External memory SERAC (Mitchell et al., 2022b)
MeLLo (Zhong et al., 2023)
In-context learning IKE (Zheng et al., 2023)
Decoding DeepEdit (Wang et al., 2024)
Table 1: Overview of current knowledge editing methods. The methods are categorized into two major families, namely parameter-modifying and parameter-preserving methods, each containing several strategies. Methods marked with $\dagger$ have the ability to process sequential edits.
2 Overview of Knowledge Editing
2.1 Problem Definition
Knowledge editing for LLMs entails modifying the output of LLMs in response to specific edit queries, with the aim of minimizing alterations to their original behavior (Yao et al., 2023a; Mazzia et al., 2023; Zhang et al., 2024a). In this section, we follow the notation from Mazzia et al. (2023).
We denote the input and output space as $\mathbb{X}$ and $\mathbb{Y}$, respectively. The function space $\mathbb{F}:\mathbb{X}\rightarrow\mathbb{Y}$ is estimated by the base model $f_{\theta_{0}}$ parameterized by $\theta_{0}\in\Theta$. Finally, let $Z_{e}=\{(x_{e},y_{e})\ |\ f_{\theta_{0}}(x_{e})\neq y_{e}\}$ be the set of edit queries we would like to apply to the base model. The goal of knowledge editing is to efficiently derive the edited model $f_{\theta_{e}}$ from the base model that satisfies the following:
$f_{\theta_{e}}(x_{e})=y_{e},\forall(x_{e},y_{e})\in Z_{e}$ (1)
The ideal edited model $f_{\theta_{e}}$ should satisfy three properties: reliability, generalization, and locality. An illustration is shown in Figure 2.
Reliability
Given an edit query $(x_{e},y_{e})$, the edited model $f_{\theta_{e}}$ should output the target answer $y_{e}$ when given the target input $x_{e}$, i.e. $f_{\theta_{e}}(x_{e})=y_{e}$. The reliability of a editing method is measured by calculating the average edit success rate:
$\mathbb{E}_{(x_{e}^{\prime},y_{e}^{\prime})\sim Z_{e}}\mathbbm{1}\{f_{\theta_{e}}(x_{e}^{\prime})=y_{e}^{\prime}\}$ (2)
Generalization
The edited model should generalize the edited knowledge to relevant instances. The generalization metric is commonly formulated as the average success rate on the neighboring set:
$\mathbb{E}_{(x_{e}^{\prime},y_{e}^{\prime})\sim N(x_{e},y_{e})}\mathbbm{1}\{f_{\theta_{e}}(x_{e}^{\prime})=y_{e}^{\prime}\},$ (3)
where $N(x_{e},y_{e})$ is the set of neighboring instances of an edit query $(x_{e},y_{e})$. Earlier works evaluate this metric by rephrasing the input prompts (Mitchell et al., 2022a; Meng et al., 2022; Huang et al., 2023).
Locality
The editing process should not affect instances unrelated to the edit queries. The locality set of an edit query $(x_{e},y_{e})$ can be defined as $L(x_{e})=\{(x_{loc},y_{loc})\in\mathbb{X}\times\mathbb{Y}\ \mathrm{s.t}\ x_{loc}\notin N(x_{e},y_{e})\land f_{\theta_{0}}(x_{loc})=y_{loc}\}$. The locality, also known as specificity, of a editing method is measured by calculating the level of invariance of model output before and after the edits, which can be calculated as follows:
$\mathbb{E}_{(x_{loc},y_{loc})\sim L(x_{e})}\mathbbm{1}\{f_{\theta_{e}}(x_{loc})=y_{loc}\}$ (4)
2.2 Current Methods
In this section, we introduce the current knowledge editing methods. The methods are categorized into parameter-modifying (Section 2.2.1) and parameter-preserving (Section 2.2.2) editing methods, each containing several strategies. An overview and illustration of current methods are included in Table 1 and Figure 3, respectively.
2.2.1 Parameter-Modifying
This category of methods, including meta-learning methods and locate-and-edit strategies, update LLMs’ knowledge by modifying their parameters.
Meta-learning
Meta-learning methods train a hyper-network to predict the update of network parameters. For instance, KnowledgeEditor (Cao et al., 2021) trains a deep network to predict weight updates. MEND (Mitchell et al., 2022a) decomposes the gradient matrix into two rank-one matrices and utilized a hyper-network to update these matrices, thereby accelerating the editing process. Built upon MEND, MALMEN (Tan et al., 2024) refines the process by formulating the aggregation of parameter shifts into a least-squares problem, further improving the scalability of meta-learning methods.
Locate and Edit
Locate-and-edit methods identify specific knowledge locations in LLMs for consequent editing. KN (Dai et al., 2022) utilizes the proposed knowledge attribution method to pinpoint neurons expressing relational facts, allowing efficient updates or erasures without fine-tuning. ROME (Meng et al., 2022) proposes the causal tracing method for identifying neuron activations associating with specific knowledge. The authors demonstrate the significance of middle-layer feed-forward networks (FFNs) in factual predictions when processing the subject’s last token. Built upon the hypothesis that the FFN modules in a transformer layer can be viewed as key-value memories (Geva et al., 2021), ROME (Meng et al., 2022) injects new knowledge into the key-value memories by deriving the closed form solution from the least-squares problem. MEMIT (Meng et al., 2023) scales up ROME (Meng et al., 2022) by editing a set of MLPs from consecutive middle-layers via solving a normal equation. PMET (Li et al., 2024a) proposes to update multi-head self-attention (MHSA) modules in addition to FFNs. EMMET (Gupta et al., 2024b) on the other hand, integrates the objectives of ROME and MEMIT into a unified preservation-memorization objective, facilitating batch-editing capabilities for both methodologies.
2.2.2 Parameter-Preserving
Parameter-preserving methods alter LLM output by adding new parameters, integrating external memory, or using strategies like in-context learning and devised decoding, without changing pre-trained LLM.
Additional Parameters
Some methods utilize additional parameters, such as adding new neurons or employing parameter-efficient techniques. CaliNET (Dong et al., 2022) extends the FFN modules with calibration memory slots to adjust the predicted token distribution. T-Patcher (Huang et al., 2023) adds neurons in the FFN’s last layer to rectify classification errors and incorrectly generated tokens, activating only in response to associated mistakes. GRACE (Hartvigsen et al., 2023) wraps a selected layer with an Adaptor that includes a codebook and deferral mechanism, learning to decode desired outputs while caching input error embeddings. The GRACE layer stores the edits and could be updated continuously over long deployments. MELO (Yu et al., 2024) utilizes DyLoRA (Valipour et al., 2023) modules to learn edits, indexing them in an inner vector database to dynamically activate corresponding LoRA blocks during inference.
External Memory
Other methods utilize external memories for editing. SERAC (Mitchell et al., 2022b) leverages a scope classifier to determine whether an user-supplied edit example stored in its memory is related to the inputs. If no example exists, the inputs are passed to the base model; otherwise, a counterfactual model generates modified answers using the inputs and the related example. MeLLo (Zhong et al., 2023) decomposes a multi-hop question into subquestions iteratively. The model then checks if the tentative answer generated by the base model contradicts the most relevant facts retrieved from the edited fact memory and adjusts the outputs accordingly.
In-Context Learning and Decoding
Certain strategies require no additional parameters. IKE (Zheng et al., 2023) edits factual knowledge via in-context learning with demonstrations to guide the language model. DeepEdit (Wang et al., 2024) employs decoding constraints, including filtering step candidates, depth-first search to store valid candidates in a stack, and a greedy search to output the optimal path for multi-hop reasoning.
3 Challenges of Knowledge Editing
While knowledge editing methods have been extensively researched, there’s a lack of comprehensive study on related challenges. In this section, we discuss the pitfalls of knowledge editing from three perspectives: inability to logically inference and robustly generalize (Section 3.1), unintended alteration of non-target knowledge (Section 3.2), and deterioration of general LLM abilities (Section 3.3).
Challenge Benchmark Metric
Portability and Generalization RippleEdits (Cohen et al., 2023) Logical Generalization
Compositionality I
Compositionality II
ConflictEdit (Li et al., 2024b) Conflict Score
Conflict Magnitude
Success Score
MQuAKE (Zhong et al., 2023) Edit-wise Success Rate
Instance-wise Accuracy
Multi-hop Accuracy
ReCoE (Hua et al., 2024) QA Accuracy
ZsRE + CounterFact Yao et al. (2023b) Subject-Replace
Reverse-Relation
One-Hop
Locality RippleEdits (Cohen et al., 2023) Subject Aliasing
Preservation
Relation Specificity
RoundEdit (Li et al., 2024b) Success Score
Distortion ($\downarrow$)
Ignore Rate ($\downarrow$)
Failure Rate ($\downarrow$)
Tied Fact Damage ($\downarrow$)
CounterFact Yao et al. (2023b) Other-Attribution
Distract-Neighbor
Other-Task
CounterFact Meng et al. (2022) Locality
Neighborhood Score
Neighborhood Magnitude
CounterFact+ Hoelscher-Obermaier et al. (2023) Neighborhood KL Divergence
Table 2: Performance benchmarks and evaluation metrics addressing generalization/portability and locality issues in knowledge editing methods. Unless specifically indicated by a downward arrow, higher values signify better performance in those evaluation metrics. CounterFact benchmark is proposed by  Meng et al. (2022), and CounterFact with mark is modified to further examine the proposed metrics.
3.1 Inability to Logically Inference and Robustly Generalize
When a fact is updated, it is crucial not only to revise the specific piece of knowledge but also to evaluate the impact on the related reasoning chain. Recently the term portability has been proposed in (Yao et al., 2023b) to evaluate the consequences after an edit, and further assess the robustness of generalization. In their study, they introduce three metrics to evaluate portability: Subject Replace (checking if synonyms of the subject are edited), Reversed Relation (checking if the reversed relation of the target is edited), and One Hop (assessing if modified knowledge is usable for further derivation). Similarly, RippleEdits benchmark as well as corresponding Logical Generalization and Compositionality metrics are proposed to examine whether edited knowledge can be inferred in composite relations of facts Cohen et al. (2023). Additionally, ReCoE benchmark is proposed to assess the propagation of updates in interconnected facts using various reasoning schemes in complex question-answering datasets Hua et al. (2024). Furthermore, MQuAKE benchmark is introduced to evaluate more complex reasoning and inference ability on multi-hop questions Zhong et al. (2023).
When editing multiple logically related facts simultaneously, models may suffer from confusion due to conflicts.  ConflictEdit benchmark is proposed to examine different editing methods on conflicted edit facts Li et al. (2024b). The different benchmarks and corresponding metrics and are arranged systematically in Table 2.
3.2 Unintended Alteration of Non-Target Knowledge
Locality is conventionally assessed using a locality dataset to evaluate edits on unrelated facts by measuring the Neighborhood Score and Neighborhood Magnitude (NS & NM; Meng et al., 2022, 2023).However, current evaluation methods do not adequately capture the post-edit effects on content beyond the locality dataset, which means the edited model could still contain unintended alterations. For example, while the location of the Louvre might be successfully changed from Paris to London, the edited model may inadvertently increase the likelihood of semantically related words (e.g. Big Ben for London) while mentioning Louvre. Some modified benchmark (CounterFact+) and corresponding metric (Neighborhood KL Divergence(Hoelscher-Obermaier et al., 2023) is then designed to disclose these previously implicit pitfalls. Another study (Yao et al., 2023a) extends this exploration to three facets of locality: Other Relations(evaluating the retention of other attributes of the updated subject), Distract Neighborhood(assessing the divergence of unrelated input when juxtaposed with edited instances), and Other Tasks (examining the influence of edits on the performance of other tasks).
Unintended edits to unrelated facts may occur because a single edit can implicitly change the predictive distribution among objects associated with the same (subject - relation) pair. After multiple consecutive edits, these alterations can accumulate and distort the stored knowledge. To evaluate this condition, the concept of Knowledge Distortion has been introduced by (Li et al., 2024b), which estimates the Jensen–Shannon divergence of the object set distribution before and after editing. This can be further extended to metrics such as the Ignore Rate, measuring how objects other than the target in the object set are neglected after editing, and the Failure Rate, which measures the proportion of instances where over half of the objects in the set are overlooked.
3.3 Deterioration of General LLM Abilities
Current evaluation metrics are primarily limited to scenarios where editing is performed only once or infrequently, prompting some studies to extend evaluations to the outcomes after consecutive edits. A study by  (Gupta et al., 2024a) discovers that post-edit models exhibit susceptibility to both gradual forgetting and catastrophic forgetting in sequential editing scenarios. Notably, their findings indicate that the extent of knowledge forgetting is more pronounced in meta-learning-based methods compared to locate-and-edit methods. Additionally, models with parameters modified successively show a decline in performance across various downstream NLP tasks (Gu et al., 2024). Furthermore, perplexity is found to increase after consecutive edits across all parameter-modified methods and different LLMs, and is proposed as another metric to indicate model collapse (Yang et al., 2024). These findings further corroborate that model editing aimed at modifying parameters adversely affects the general capabilities of the original LLMs.
Single Edit Multiple Edit
One-Hop Multiple-Hop Reverse Conflict Composite Conflict
Methods SR RR OH MH CS CM CS CM
FT 72.96 8.05 1.34 1.6 80.28 71.11 75.45 64.28
MEND 42.45 0.00 11.34 9.2 88.89 60.50 84.85 43.45
ROME 37.42 46.42 50.91 7.6 65.92 -0.65 71.70 37.04
MEMIT 27.73 47.67 52.74 8.1 51.40 -1.60 57.15 -1.50
SERAC 17.79 1.30 5.53 7.9 50.89 -0.02 50.84 -0.02
IKE 88.77 92.96 55.38 8.3 58.20 -1.00 50.52 -0.99
Table 3: Experimental results for portability and generalization. SR: Subject-Replace, RR: Reverse-Relation, OH: One-Hop Accuracy, EW: Edit-wise, IW: Instance-wise, MH: Multi-hop Accuracy, CS: Conflict score, CM: Conflict magnitude. Higher values indicate better performance for all metrics in this table. Results marked with $\dagger$ are obtained in our own experiments, and other results are taken from previous studies.
Single Edit Multiple Edit
Methods OA DN OT Succ. D ($\downarrow$) IR ($\downarrow$) FR ($\downarrow$)
FT 12.88 9.48 49.56 100.0 16.12 97.48 97.32
MEND 73.50 32.96 48.86 99.12 14.35 87.64 86.56
ROME 78.94 50.35 52.12 99.80 13.95 78.98 77.60
MEMIT 86.78 60.47 74.62 99.72 13.50 72.03 70.44
SERAC 99.50 39.18 74.84 50.14 3.78 99.62 99.64
IKE 84.13 66.04 75.33 100.0 13.43 73.53 73.00
Table 4: Experimental results for locality. OA: Other-Attribution, DN: Distract-Neighbor, OT: Other-Task, Succ.: Success rate, D: Distortion, IR: Ignore rate, FR: Failure rate, TFD: Tied fact damage. Unless specifically indicated by a downward arrow, higher values signify better performance in those evaluation metrics. Results marked with $\dagger$ are obtained in our own experiments, and other results are taken from previous studies.
4 Experiments
The experiments were done to evaluate robust generalization and locality (Section  4.1.1 as well as deterioration of general LLM abilities (Section  4.1.2 across different editing methods.
4.1 Experimental Setup
4.1.1 Robust generalization and locality
We use GPT-J (Wang and Komatsuzaki, 2021) as the baseline model for editing and implement six distinct editing methodologies to assess robust generalization and locality: MEND (meta-learning), ROME and MEMIT (locate-and-edit), SERAC (external memory), and IKE (prompting demonstrations).
Given the overlap in benchmarks for robust generalization and locality, we select a subset for our experiments. Robust generalization is evaluated under single edit and multiple edit settings. Single edit metrics include Subject-Replace, Reverse-Replace, and One-Hop reasoning (Yao et al., 2023a). Multiple edit metrics include multi-hop editing accuracy (Zhong et al., 2023), and Conflict Score and Conflict Magnitude for Reverse Conflict and Composite Conflict respectively (Li et al., 2024b).
For locality, single edit metrics include Other-Attribution, Distract-Neighbor, and Other-Task (Yao et al., 2023b), while multiple edit metrics encompass Success Rate, Distortion, Ignore Rate, and Failure Rate (Li et al., 2024b).
4.1.2 Deterioration of general LLM abilities
Following the settings of (Gu et al., 2024), we assessed deterioration of general LLM abilities post-editing using six methodologies: ROME, MEMIT, SERAC, MEND, KN, and GRACE. We evaluated general abilities across four NLP downstream tasks: open-domain question answering, sentiment analysis, reasoning, and summarization. These tasks were assessed after 10 to 40 edits on the Zero-Shot Relation Extraction (ZsRE) dataset(Levy et al., 2017), comparing the results against pre-editing benchmarks. More details on the selected downstream tasks are in Appendix B.
4.2 Experimental Results and Discussion
In general, current editing methodologies exhibit suboptimal performance concerning both robust generalization and locality. Regarding robust generalization, IKE, which leverages prompt demonstrations, demonstrates superior performance over other methodologies in single edit conditions. However, IKE’s performance noticeably declines in multiple edit scenarios, suggesting that prompt demonstrations may become confused when editing multiple logically related facts. Conversely, fine-tuning and meta-learning-based methods are less susceptible to confusion after editing multiple related facts.
Regarding locality, IKE maintains stable performance across metrics in single edit settings. Parameter-modifying methods excel in Other Attribution but decline in other metrics, except MEMIT, which remains stable across all metrics. In multiple edit scenarios, all methods except SERAC show similar performance. In the multiple edit scenario, all methods except SERAC exhibit relatively similar performance. SERAC displays low edit success rate and distortion rate, suggesting its scope classifier does not adopt most edits in this scenario. This may be attributed to its weakness in recovering edited facts, which is crucial in this metric setting.
In terms of general LLM abilities, the number of edits affects methods differently. Meta-learning methods like MEND degrade significantly after 10-20 edits. Locate-and-edit methods such as ROME and KN degrade after 10 edits, while MEMIT remains stable after 40 edits. This disparity can be attributed to MEMIT’s strategy of adjusting parameters across multiple layers, as opposed to ROME’s single-layer edits and KN’s approach of modifying a few neurons. This distribution of parameter modifications across layers helps mitigate deterioration.
GRACE, which stores edited facts with additional parameters, shows no performance change in downstream tasks after edits. One possible explanation is that the edits are conducted on the ZsRE dataset, which is distinct from the requirements of downstream tasks, leading to the stored facts not being retrieved during inference. Similarly, SERAC, utilizing external memory for edited facts, preserves general NLP abilities post-editing. This preservation stems from SERAC being trained once before editing begins, solely performing inference during editing, thereby preventing changes in the model’s output, even after multiple edits.
Overall, parameter-modifying methods degrade downstream task performance by altering pre-trained LLM parameters. In contrast, parameter-preserving methods maintain the original parameters, resulting in stable downstream task performance even after multiple edits.
(a) Open Domain Question Answering
(b) Sentiment Analysis
(c) Summarization
(d) Reasoning
Figure 4: The experimental results for the deterioration of general abilities were obtained by editing a pretrained LLM using GPT-J with various editing algorithms, including ROME, MEMIT, MEND, KN, SERAC, and GRACE, each applied 10 to 40 times. The edited models were subsequently evaluated on four downstream tasks, including open-domain question answering, sentiment analysis, summarization, and reasoning.
5 Future Prospects
5.1 Leveraging Information Retrieval and External Memory
Previous research has demonstrated the benefits of utilizing external knowledge bases, as opposed to relying solely on internal knowledge, to guide LLMs in generating content based on a predefined set of facts. These methods effectively separate the factual knowledge stored in LLMs from the inference processes, thus reducing potential biases encoded within the models.
External knowledge bases can include diverse sources such as extensive text corpora, structured tables, or even simple key-value databases. Once these knowledge sources are provided, one can either finetune the LLMs to enhance their ability to retrieve information or employ prompting and in-context learning techniques to query these sources while keeping the model parameters intact. Such approaches not only eliminate the need to verify and edit false factual knowledge within the LLMs but also facilitate the use of attribution and reflection methods. This ensures that the generated content aligns with the predefined external knowledge base, thereby enhancing both accuracy and accountability.
5.2 Improve Understandings of LLMs’ Internal Knowledge Structures
While the identification of factual knowledge storage in LLMs has been extensively explored in recent literature (Meng et al., 2022, 2023; Dai et al., 2022; Hernandez et al., 2024; Geva et al., 2021), the correlation between the location of knowledge and the success rate of model editing remains low (Hase et al., 2023). Additionally, despite evidence suggesting a strong connection between factual knowledge and the feed-forward network layers (Meng et al., 2022; Geva et al., 2021, 2022), recent findings (Li et al., 2024a) indicate that updates to multi-head self-attention layers also lead to improved outcomes. These studies highlight that merely locating fact storage does not fully elucidate the underlying mechanisms of knowledge structures in LLMs. Therefore, further research into how knowledge locations interact with model predictions is essential for advancing the interpretability and controllability of LLMs.
In addition to enhancing the success rate of edits, preserving the general capabilities of LLMs is crucial for assessing the efficacy of model editing methods, as discussed in Section 3.3. Recent breakthroughs in identifying regions within models that correlate with general linguistic abilities (Zhang et al., 2024b) have opened up a direction for future research in model editing. Specifically, by locating these critical areas, it is possible to perform targeted modifications while keeping these regions away from alterations, thereby preventing the deterioration of general abilities. Consequently, advancements in related fields would ensure that edits could be performed without compromising the overall performance of the LLMs, thereby significantly enhancing the specificity and effectiveness of current model editing methods.
5.3 Improve Robustness of Knowledge Editing
Even after successful (achieving fair scores on the existing metrics) edit, the revised model may refuse the modification if the knowledge concerning the altered concept is challenged by extended dialogues. Instead, it might revert to the pre-edit version (reversion), or output an ambiguous answer about the edited concept (confusion). Given the interconnected nature of knowledge, experiments disclose that more popular the knowledge in the benchmark, the easier for the modified model to trace back the original concept (Ma et al., 2024). It highlights the unsatisfying robustness of the existing editing strategies. A more comprehensive understanding how LLMs store and process among different knowledge entities is crucial for a more robust editing. We are also short of specific benchmarks and automated metrics addressing on theses issues. Knowledge-focused editing would not avoid the hallucination inherited from the pre-edit model. TruthX (zhang2024truthx) tries to alleviate hallucination via a parameter-preserved approach by mapping the LLMs internal representation to semantic and truthful spaces and edits the truthfulness in the truthful space. Combination of truthfulness and knowledge adjustment in the same space may provide a practical solution.
6 Conclusion
Although model editing techniques appear promising for cost-effectively updating knowledge, they still have significant pitfalls. Current editing methods often struggle with making logical inferences based on the edited facts, introducing unintended alterations of non-target knowledge and deterioration in model performance, particularly with parameter-modified methods. Editing techniques that leverage information retrieval can mitigate deviations in model abilities by keeping model parameters intact, as demonstrated in our experiments. Moreover, gaining a deeper understanding of how models store and process knowledge can enhance the controllability of edited facts, leading to greater robustness. We hope our work illuminates potential directions for future improvements in knowledge editing.
Appendix A Detailed Explanation of Evaluation Metrics and Examples
A.1 Portability / Generalization
Single Edit
In the single edit scenario, we further classify the methods into two settings:
  • One-Hop: This setting focuses on evaluating the impact of a single edit on direct, one-hop reasoning tasks.
  • Multi-Hop: This setting assesses the impact of a single edit on more complex, multi-hop reasoning tasks.
For one-hop evaluations, we adopt the methods proposed by (Yao et al., 2023a). These include:
  • Subject Replace: This metric tests the model’s generalization ability by replacing the subject in the question with an alias or synonym, assessing if the edited attribute is generalized to other descriptions of the same subject.
  • Reversed Relation: This metric evaluates the model’s capability to handle reversed relations by filtering for suitable relations such as one-to-one and asking the reverse question to check if the target entity is also updated.
  • One-Hop Test: This metric assesses the edited language model’s performance on downstream tasks that require one-hop reasoning.
In the multi-hop setting, we assess the model’s performance on multi-hop questions using the evaluation methods proposed by (Zhong et al., 2023), which include:
  • Edit-wise Success Rate (EW): This metric measures how many facts can be successfully recalled from the edited language model.
    $\mathrm{EW}=\mathbbm{1}\{f^{*}(s)=o^{*}\}$ (5)
    where $f^{*}$ is the model after editing.
  • Instance-wise Accuracy (IW): This metric tests how many multi-hop instances the model can recall all the individual single-hop facts. This metric is crucial for multi-hop performance, as the model must encode each fact to answer the multi-hop question.
    $\mathrm{IW}=\mathbbm{1}\{\bigwedge_{(s,r,o^{*})\in C^{*}}[f^{*}(s)=o^{*}]\}$ (6)
    where $C^{*}=\langle(s_{1},r_{1},o_{1}),\dots,(s_{n},r_{n},o_{n})\rangle$ is the chain of facts of a multi-hop question. In this chain, the object of the $i^{\mathrm{th}}$ fact is the subject of the next fact. (i.e., $o_{i}=s_{i+1}$)
  • Multi-hop Accuracy (MH): This metric assesses the accuracy of the original and edited language models on multi-hop questions. In the MQuAKE dataset (Zhong et al., 2023), there are three generated multi-hop questions for each instance. If any of the three questions is correctly answered by the model, we consider it accurate.
    $\mathrm{MH}=\mathbbm{1}\{\bigvee_{q\in Q}f^{*}(q)=a^{*}\}$ (7)
    where $Q$ is a set of similar multi-hop questions with the same answer $a^{*}$.
Multiple Edits
In the multiple edits scenario, we test the model’s performance after applying multiple edits. For this, we use the setting and evaluation methods from (Li et al., 2024b). The settings consist of:
  • Reverse Conflict: This setting introduces conflicts by editing facts with reverse relations. For example:
    edit 1: ($s_{1},r_{1},o_{1}$$o_{2}$)
    Hamlet was written by ShakespeareAgatha Christie.
    edit 2: ($o_{2},r_{2},s_{1}$$s_{2}$)
    The notable work of Agatha Christie is HamletOdyssey
    the updated knowledge then could be represented as:
    $\left\{\begin{array}[]{l}k_{0}=(s_{1},r_{1},o_{2})\\ k_{n}=(s_{2},r_{1},o_{2})\end{array}\right.$
  • Composite Conflict: This explores more complex situations where the edits are associated with a fact that is not influenced by the editing (tied fact). For example:
    edit 1: ($s_{1},r_{1},o_{1}$$o_{2}$)
    Hamlet was written in EnglishFrench
    edit 2: ($s_{2},r_{2},o_{2}$$o_{3}$)
    Shakespeare wrote in FrenchGerman
    tied fact: ($s_{1},r,s_{2}$)
    The notable work of Shakespeare is Hamlet
    where $r\land r_{1}\to r_{2}$ is a logical rule. The updated knowledge then could be represented as:
    $\left\{\begin{array}[]{l}k_{f}=(s_{1},r,s_{2})\\ k_{0}=(s_{1},r_{1},o_{2})\\ k_{n}=(s_{1},r_{1},o_{3})\end{array}\right.$
The evaluation methods include:
  • Conflict Score (CS): Measures how well a knowledge editing method handles knowledge conflicts by calculating the ratio that the new fact is more probable than the old fact after knowledge editing.
    $\mathrm{CS=}\mathbbm{1}\{p_{f_{\theta}^{\prime}}(k_{n})>p_{f_{\theta}^{\prime}}(k_{o})\}$ (8)
  • Conflict Magnitude (CM): Estimates the decrease in probability of the old fact after editing.
    $\mathrm{CM=}\frac{p_{f_{\theta^{m}}}(k_{o})-p_{f_{\theta^{\prime}}}(k_{o})}{p_{f_{\theta^{m}}}(k_{o})}$ (9)
    $\theta^{m}$ is the intermediate model parameters after edit 1.
A.2 Locality
Single Edit
In the single edit scenario for locality, we adopt the methods proposed by (Yao et al., 2023b), including:
  • Other Attribution The modified CounterFact dataset is applied to test whether the non-target attributes of the edited subjects remained the same. For example, if we reset Lionel Messi as a basketball player, his nationality should stay the same.
  • Distract Neighbor Pretend model editing by modifying the neighborhood prompt in the CounterFact. For example, if the original prompt is "Windows 11 is a product of __", the modified prompt would be "Windows 11 is a product of Google. Office 365, developed by __". It testifies whether the model prediction would be "distracted" by the revised prompt.
  • Other Task The edited model is tested on the multiple-choice QA task Physical Interaction QA(PIQA, Bisk et al. (2020)) and the performance is evaluated by accuracy.
Multiple Edits
We also test the model’s locality in the multiple edits scenario adopting the methods and evaluations from (Li et al., 2024b). The settings consist of:
  • Round Edit: This edits the knowledge triplet back-and-forth, for example:
    edit 1: ($s,r,o_{1}$$o*$)
    edit 2: ($s,r,o*$$o_{1}$)
the evaluation metrics include:
  • Distortion (D) (Li et al., 2024b):
    $D=JS\left(p_{f_{\theta}}(\text{Obj}\mid(s,r)),p_{f_{\theta^{\prime}}}(\text{Obj}\mid(s,r))\right)$ (10)
    estimates the JS divergence of the objects distribution before and after edit.
  • Ignore Rate (IR) (Li et al., 2024b):
    $\begin{split}\mathrm{IR=}\frac{1}{\left|\mathrm{Obj}\right|-1}\sum_{\begin{subarray}{c}o\in\mathrm{Obj}\setminus\{o1\}\end{subarray}}&\mathbbm{1}\{p_{f_{\theta}}(o\mid(s,r))>\\ &p_{f_{\theta}^{\prime}}(o\mid(s,r))\}\end{split}$ (11)
  • Failure Rate (FR) (Li et al., 2024b):
    $\mathrm{FR=}\mathbbm{1}\{\mathrm{IR}>0.5\}$ (12)
  • Tied Fact Damage (TDF) (Li et al., 2024b):
    $\mathrm{TFD=}\frac{p_{f_{\theta^{m}}}(k_{f})-p_{f_{\theta^{\prime}}}(k_{f})}{p_{f_{\theta^{m}}}(k_{f})}$ (13)
    $k_{f}$ denotes the tied facts and $\theta^{m}$ is the intermediate model parameters after edit 1.
Other Locality Metrics
  • Neighborhood KL Divergence (Hoelscher-Obermaier et al., 2023):
    $\mathrm{NKL}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{w\in W}\log\left(\frac{P(w)}{P^{*}(w)}\right)$ (14)
  • Neighborhood Score (NS) (Meng et al., 2022): collect a set of "neighbor" subjects and evaluation the success fraction for $\mathbb{P}\left[o^{c}\right]>\mathbb{P}\left[o^{*}\right]$, while the $o^{c}$ denotes the correct facts and $o^{*}$ denotes the false facts.
  • Neighborhood Magnitude (NM) (Meng et al., 2022): the differences of $\mathbb{P}\left[o^{c}\right]$ and $\mathbb{P}\left[o^{*}\right]$ for the "neighborhood" subjects.
Appendix B Detailed Explanation of experiments for deterioration of general LLM abilities
We follow the settings of  (Gu et al., 2024) for this part of experiments. Different evaluation metrics were applied for each downstream task: Exact Match for open-domain question answering on the Natural Question dataset  (Kwiatkowski et al., 2019), accuracy for sentiment analysis on the SST2 dataset  (Socher et al., 2013), solve rate for reasoning on the GSM8K dataset  (Cobbe et al., 2021), and ROUGE score for summarization on the SAMSum dataset  (Wang et al., 2019).